OpenStudy (anonymous):

Will Give For Best Explanation And Answer! Simplify the radical expression by rationalizing the denominator. 2sqrt108 / sqrt180y

OpenStudy (softballgirl372015):

hi! I solved the problem and got the answer \[\frac{ 2\sqrt{15y} }{ 5y}\] Do you need me to explain it?

OpenStudy (anonymous):

yes, that would be very helpfull

OpenStudy (softballgirl372015):

okay! So what I did first was used prime factorization to get the radicals to its prime numbers. That gave me the fraction of \[\frac{ 12\sqrt{3} }{ 6\sqrt{5y}}\] Then I multipled the numerator and denominator by \[\sqrt{5y}\] so that the radical would be eliminated from the denominator. When I multipled the numerator and denominator, I got \[\frac{ 12\sqrt{3*5y} }{ 6*5y }\] If you simply this, you then get \[\frac{ 12\sqrt{15y} }{ 30y}\] which can be further simplified to \[\frac{ 2\sqrt{15y} }{ 5y }\] which is your final answer.

OpenStudy (anonymous):

Wow, your a genius! Thank you so much!

OpenStudy (softballgirl372015):

No problem :)